# Lower-mass-gap Black Holes in Dense Star Clusters

### v1

@@ Abstract
The existence of compact stellar remnants in the mass range has long been debated. This so-called lower mass gap was initially suggested by the lack of low-mass X-ray binary observations with accretors in this mass range, but it has recently been called into question following newer observations, including a lower-mass-gap candidate with a millisecond pulsar companion in the dense globular cluster NGCĀ 1851. Here we model NGCĀ 1851 with a grid of similar dense star clusters utilizing the state-of-the-art Monte Carlo -body code CMC, and we specifically study the formation of lower-mass-gap black holes. We demonstrate that both massive star evolution and dynamical interactions can contribute to forming lower-mass-gap black holes. In general, the collapse of massive remnants formed through mergers of neutron stars or massive white dwarfs produces the largest number of lower-mass-gap black holes among all formation channels. However, in more massive clusters, supernova core collapse can contribute comparable numbers. Our NGCĀ 1851-like models can reproduce millisecond pulsar ā lower-mass-gap black hole binaries similar to the observed system. Additionally, the lower-mass-gap black holes can also become components of dynamically assembled binaries, and some will be in merging black holeāneutron star systems similar to the recently detected gravitational wave source GW230529. However, the corresponding merger rate is probably .
@@ 1 Introduction
The maximum mass of neutron stars (NSs) and the transitional mass between NSs and black holes (BHs) have been long-standing questions in astrophysics. Theoretical calculations show NSs have gravitational masses (e.g., Rhoades & Ruffini, 1974; Kalogera & Baym, 1996), while pulsar timing observations have identified (Cromartie etĀ al., 2020; Fonseca etĀ al., 2021) as the most massive NS to date. At the same time, dynamical mass measurements of BHs in X-ray binaries have suggested a lack of BHs with mass , indicating the possible existence of a mass gap, usually referred to as the ālower mass gapā (henceforth LMG), between and (Bailyn etĀ al., 1998; Ćzel etĀ al., 2010; Farr etĀ al., 2011). These findings inspired searches for the origins of the LMG, including those resulting from core-collapse supernovae (Fryer etĀ al., 2012; Belczynski etĀ al., 2012; Kochanek, 2014; Fryer etĀ al., 2022) and binary evolution of massive stars (e.g., Siegel etĀ al., 2023).
@@
Recently, yet another LMG candidate was identified in the Galactic globular cluster NGCĀ 1851 (Barr etĀ al., 2024). This object, with inferred mass (95% confidence interval), was found via radio timing of its companion, the millisecond pulsar PSR J0514ā4002E (NGCĀ 1851E). This could be the most massive NS or least massive BH detected in a millisecond pulsar (MSP) binary to date. MSPs are abundant in globular clusters (GCs) and are understood to form at high rates in these systems due to the role of dynamical processes enabled by high stellar densities. As pointed out by Barr etĀ al. (2024), the companion of NGCĀ 1851E bears a striking resemblance in mass to the remnant of the known double neutron star binary merger, GW170817, observed by LIGO-Virgo-KAGRA (LVK; Abbott etĀ al., 2017), potentially hinting that this object formed via a similar mechanism. The double NS merger origin is also motivated by the presence of a known double NS binary with an inspiral time of roughly Myr in another Galactic globular cluster, M15 (Jacoby etĀ al., 2006). Additional formation scenarios are also viable including mergers of massive white dwarf binaries and accretion-induced collapse via binary Roche lobe overflow or tidal disruption of stars (e.g., Kremer etĀ al., 2021, 2022; Ye etĀ al., 2024). Furthermore, āless exoticā scenarios involving single-star evolution cannot be ruled out due to the uncertainties in supernova natal kicks (e.g., OāDoherty etĀ al., 2023, and references therein). However, retention within a GC with central escape velocity places constraints on the natal kick such an object could have received at birth (e.g., Ye etĀ al., 2019, and references therein).
@@
In this study, we explore the formation and dynamics of LMG objects in dense stellar clusters, similar to the companion of NGCĀ 1851E. We focus in particular upon the formation of these objects via NS collapse, triggered by several mechanisms enabled by dynamical interactions including double NS mergers, white dwarf mergers, and stellar tidal disruptions. We then explore the subsequent dynamical evolution of these mass-gap objects in clusters including their interactions with BHs, NSs, and other stars. We demonstrate that objects like the NGCĀ 1851E binary can form naturally in clusters like NGCĀ 1851 and predict other related GW/electromagnetic sources that may potentially be observed in the future.
@@ 2 Monte Carlo -body Code and Compact Object Formation
We use the Cluster Monte Carlo code (CMC; Rodriguez et al., 2022, and references therein) to study the formation of LMG BHs in GCs. CMC is a Monte Carlo -body code based on the orbit-averaged method of Hénon (1971a, b). It incorporates various relevant physics for cluster evolution, including two-body relaxation, strong three- and four-body dynamical interactions, and tidal mass loss. The Fewbody package is used to integrate directly strong three- and four-body gravitational encounters (Fregeau et al., 2004; Fregeau & Rasio, 2007), which includes post-Newtonian dynamics for BHs (Antognini et al., 2014; Amaro-Seoane & Chen, 2016; Rodriguez et al., 2018a, b). Single and binary star evolution is fully coupled to the stellar dynamics and is computed with COSMIC (Breivik et al., 2020), which is based on SSE (Hurley et al., 2000) and BSE (Hurley et al., 2002).
@@ 2.1 Core-collapse Supernova
BHs and NSs can form in core-collapse supernovae in CMC. The birth mass distribution of compact objects depends on the supernova mechanism and the formation of LMG objects may be determined by effects including the explosion energy and the development timescale of instabilities (e.g., Fryer etĀ al., 2012; Belczynski etĀ al., 2012; Fryer etĀ al., 2022), or the successfulness of the supernova explosion (e.g., Kochanek, 2014). We follow the prescriptions in Fryer etĀ al. (2012) for the remnant mass and natal kick distributions of core-collapse supernovae and adopt either their ārapid modelā or ādelayed modelā in the simulations. The delayed mechanism launches an explosion in a longer timescale than the rapid mechanism and allows for a broader range of explosion energy and fallback material onto the proto-NS. This mechanism thus produces a continuous range of remnant masses across the LMG (assuming to be between and ), in contrast to the rapid mechanism which leads to a sharp transition between BHs and NSs and produces very few remnants in the mass gap (FigureĀ 4 in Fryer etĀ al. (2012) and also the ārapidā and ādelayedā histograms in FigureĀ 2 below). We adopt a maximum NS mass at .
We assume that newly born NSs from core-collapse supernovae receive large natal kicks drawn from a Maxwellian distribution with standard deviation (Hobbs etĀ al., 2005). On the other hand, the natal kicks of BHs are determined by how much material falls back post-explosion and are calculated by
@@ 2.2 Electron-capture Supernova and White Dwarf Collapse
In addition to core-collapse supernovae, NSs can form from electron-capture supernovae in our simulations (Nomoto, 1984, 1987). Electron-capture can also be triggered in massive white dwarfs (WDs) through accretion onto the WDs or mergers between two WDs (Nomoto & Kondo, 1991; Saio & Nomoto, 1985, 1998, 2004; Shen etĀ al., 2012; Schwab etĀ al., 2015, 2016; Schwab, 2021). We assume that all super-Chandrasekhar mergers where the components are CO or ONe WDs collapse to NSs (e.g., Kremer etĀ al., 2023; Ye etĀ al., 2024), thus maximizing the NS formation rate through this channel. For the natal kicks of electron capture supernovae, we adopt the prescription in Kiel etĀ al. (2008) and Kiel & Hurley (2009) and sample the kicks from a Maxwellian distribution with standard deviation . Therefore, many NSs formed in electron capture supernovae are retained in typical GCs with escape velocities of .
@@ 2.3 Collapse of Neutron Stars
To briefly summarize, NS-main sequence TDEs occur when the pericenter distance between the two objects during a close encounter is smaller than the tidal disruption radius of the main-sequence star:
@@
where , , and are the initial mass and radius of the accretion disk, and the viscous accretion time, respectively. The exponent is a free parameter determining the amount of mass transported from the outer edge of the accretion disk to the NS (e.g., Blandford & Begelman, 1999), and (e.g., Kumar etĀ al., 2008). We assume that following the results of the hydrodynamic simulations of Kremer etĀ al. (2022). This is a reasonable assumption since the relative velocity of the NSs and main-sequence stars in typical GCs is, in general, much smaller than the stellar escape velocity. We also assume that , day, and equals the NS radius (Kremer etĀ al., 2022). For more details of the TDE treatments in CMC, see Ye etĀ al. (2024).
@@ 3.1 Typical Core-collapse Globular Clusters
We run 13 GC simulations up to a Hubble time with the same initial conditions using CMC to explore the formation of LMG BHs through core-collapse supernovae, merger-induced collapse of WDs, and tidal disruption of stars by NSs. These cluster simulations have initial number of stars , initial virial radius pc, metallicity , and Galactocentric distance kpc. We adopt the Kroupa standard broken power-law (Kroupa, 2001) for the initial mass function between and . The initial binary fraction is assumed to be . These simulations will evolve to be ātypicalā core-collapsed clusters similar to NGCĀ 6752 in the Milky Way (for more details, see Ye etĀ al., 2024). We show the surface brightness profiles and velocity dispersion profiles of all 13 simulations at Gyr in FigureĀ 1 (left panels). We focus on core-collapsed GCs in this study since they produce binaries containing NSs, low-mass BHs, or LMG objects at higher rates than non-core-collapsed GCs (e.g., Ye etĀ al., 2020). For these simulations, we vary the supernova prescriptions, the mass scenarios for WDāWD mergers, and the accretion efficiency of NSs tidally disrupting main-sequence stars (SectionĀ 2). These variations will affect the mass distributions of NSs and the number of LMG BHs formed. We also vary the prescriptions for the tidal capture of two WDs, and the stability criterion of binary mass transfer, . This stability criterion will influence the orbital period distribution of the compact object binaries, thereby affecting their cross sections for dynamical encounters (for more details, see Ye etĀ al., 2024). TableĀ 3.2 lists the different prescriptions adopted for each simulation.
To explore the extreme upper limit of the formation of LMG BHs, we also run a simulation (simulationĀ 14) which truncates the standard Kroupa initial mass function at , resulting in a much smaller number of BHs with masses greater than about . Additionally, we assume that NSs born from core-collapse supernovae receive small natal kicks same as those from electron-capture supernovae. These assumptions allow for a large number of NSs and LMG objects to be retained in the cluster. Other initial conditions of this simulation are the same as mentioned above.
@@ 3.2 Modeling NGCĀ 1851
In addition to the main grid of models which are intended to be representative of a ātypicalā core-collapsed Milky Way GC, we also run two additional higher mass simulations designed specifically to match the observed properties of NGCĀ 1851. For these models, we assume an initial number of stars and virial radius Ā pc. The metallicity of the simulations is , closely matching the observed metallicity of NGCĀ 1851 from Harris (1996, 2010 edition). We adopt a circular orbit with Galactocentric distance Ā kpc for the cluster following https://people.smp.uq.edu.au/HolgerBaumgardt/globular/, which shows that NGCĀ 1851 is on an eccentric orbit in the Milky Way with pericenter distance Ā kpc and apocenter distance Ā kpc. At distances near the pericenter distance, NGCĀ 1851 may be subject to strong tidal stripping from the Milky Way, and a simple circular orbit at the apocenter distance would not be able to capture the detailed evolution of the cluster in the Milky Wayās potential. However, simulations with Ā kpc match the observed surface brightness profile of NGCĀ 1851 at the outermost part better than those with smaller Galactocentric distances, as we will demonstrate below. This is not surprising since the cluster spends most of its orbital time near the apocenter distance, while only a short time at the pericenter distance. In addition, back scatteringāgravitational interactions with other stars scattering the potential escapers back down to a bound orbitāmay also prevent a significant fraction of objects from escaping from their host clusters even though they have enough energy to do so (e.g., Weatherford etĀ al., 2023). Simulating the detailed evolution of the clusterās tidal boundary is beyond the scope of this paper, and the location of the tidal boundary likely does not have a significant effect on the LMG objects, which mostly concentrate at the cluster cores due to mass segregation. Other initial cluster properties (e.g., binary fraction, initial mass function, stellar distribution profile) are the same as the simulations in the main grid (simulations 1-13 above). For these two simulations, we again adopt different prescriptions that will affect the mass of the NSs and BHs for comparison (TableĀ 3.2). We evolve all simulations to Ā Gyr.
@@
3.80 | 24 | 970 | 22 | 1 | 1 | delayed | updated | 0.2 | ā |
Note. ā From left to right, the columns are the total mass of the cluster at present, the number of BHs retained at present, the number of NSs retained at present, and the number of compact object binaries (NSāWD, NSāNS, and BHāNS) at present. The present-day values for simulations 1-14 are averages between 11 and 13.8Ā Gyr in each cluster simulation, and between 10 and 13Ā Gyr for simulations A-B.
@@
TableĀ 3.2 summarizes the present-day properties of the 14+2 GC simulations and their prescriptions for forming LMG BHs. The present-day values are averaged between 11 and 13.8Ā Gyr for simulations 1-14, while between 10 and 13Ā Gyr for simulations A-B (the observed age of NGCĀ 1851 is Ā Gyr; VandenBerg etĀ al., 2013). All clusters have reached core-collapsed at late times, which is shown by the small number of BHs retained (e.g., Kremer etĀ al., 2020a, and references therein). In the following sections, we will explore the contributions from various pathways to the formation of binaries and transient sources containing LMG objects.

### v2

@@ Abstract
The existence of compact stellar remnants in the mass range has long been debated. This so-called lower mass gap was initially suggested by the lack of low-mass X-ray binary observations with accretors about , but it has recently been called into question following newer observations, including a lower-mass-gap candidate with a millisecond pulsar companion in the dense globular cluster NGCĀ 1851. Here we model NGCĀ 1851 with a grid of similar dense star clusters utilizing the state-of-the-art Monte Carlo -body code CMC, and we specifically study the formation of lower-mass-gap black holes. We demonstrate that both massive star evolution and dynamical interactions can contribute to forming lower-mass-gap black holes. In general, the collapse of massive remnants formed through mergers of neutron stars or massive white dwarfs produces the largest number of lower-mass-gap black holes among all formation channels. However, in more massive clusters, supernova core collapse can contribute comparable numbers. Our NGCĀ 1851-like models can reproduce millisecond pulsar ā lower-mass-gap black hole binaries similar to the observed system. Additionally, the lower-mass-gap black holes can also become components of dynamically assembled binaries, and some will be in merging black holeāneutron star systems similar to the recently detected gravitational wave source GW230529. However, the corresponding merger rate is probably .
@@ 1 Introduction
The maximum mass of neutron stars (NSs) and the transitional mass between NSs and black holes (BHs) have been long-standing questions in astrophysics. Theoretical calculations indicate that NSs have gravitational masses (e.g., Rhoades & Ruffini, 1974; Kalogera & Baym, 1996), while pulsar timing observations have identified (Cromartie etĀ al., 2020; Fonseca etĀ al., 2021) as the most massive NS to date. At the same time, dynamical mass measurements of BHs in X-ray binaries have suggested a lack of BHs with mass , indicating the possible existence of a mass gap, usually referred to as the ālower mass gapā (henceforth LMG), between and (Bailyn etĀ al., 1998; Ćzel etĀ al., 2010; Farr etĀ al., 2011). These findings inspired searches for the origins of the LMG, including those resulting from core-collapse supernovae (Fryer etĀ al., 2012; Belczynski etĀ al., 2012; Kochanek, 2014; Liu etĀ al., 2021; Fryer etĀ al., 2022) and binary evolution of massive stars and compact objects (e.g., Gao etĀ al., 2022; Siegel etĀ al., 2023).
@@
Recently, yet another LMG candidate was identified in the Galactic globular cluster NGCĀ 1851 (Barr etĀ al., 2024). This object, with inferred mass (95% confidence interval), was found via radio timing of its companion, the millisecond pulsar PSR J0514ā4002E (NGCĀ 1851E). This could be the most massive NS or least massive BH detected in a millisecond pulsar (MSP) binary to date. MSPs are abundant in globular clusters (GCs) and are understood to form at high rates in these systems due to the role of dynamical processes enabled by high stellar densities. As pointed out by Barr etĀ al. (2024), the companion of NGCĀ 1851E bears a striking resemblance in mass to the remnant of the known double neutron star binary merger, GW170817, observed by LIGO-Virgo-KAGRA (LVK; Abbott etĀ al., 2017), potentially hinting that this object formed via a similar mechanism. The double NS merger origin is also motivated by the presence of a known double NS binary with an inspiral time of roughly Myr in another Galactic globular cluster, M15 (Jacoby etĀ al., 2006). Additional formation scenarios are also viable including mergers of massive white dwarf binaries and accretion-induced collapse via binary Roche lobe overflow or tidal disruption of stars (e.g., Kremer etĀ al., 2021, 2022; Ye etĀ al., 2024). Furthermore, scenarios involving single-star evolution cannot be ruled out due to the uncertainties in supernova natal kicks (e.g., OāDoherty etĀ al., 2023, and references therein). However, retention within a GC with central escape velocity places constraints on the natal kick such an object could have received at birth (e.g., Ye etĀ al., 2019, and references therein).
@@
In this study, we explore the formation and dynamics of LMG BHs in dense stellar clusters, similar to the companion of NGCĀ 1851E. We focus in particular upon the formation of these objects via NS collapse, triggered by several mechanisms enabled by dynamical interactions including double NS mergers, white dwarf mergers, and stellar tidal disruptions. We then explore the subsequent dynamical evolution of these mass-gap objects in clusters including their interactions with BHs, NSs, and other stars. We demonstrate that objects like the NGCĀ 1851E binary can form naturally in clusters like NGCĀ 1851 and predict other related GW/electromagnetic sources that may potentially be observed in the future.
@@ 2 Monte Carlo -body Code and Compact Object Formation
We use the Cluster Monte Carlo code (CMC; Rodriguez et al., 2022, and references therein) to study the formation of LMG BHs in GCs. CMC is a Monte Carlo -body code based on the orbit-averaged method of Hénon (1971a, b). It incorporates various relevant physics for cluster evolution, including two-body relaxation, strong three- and four-body dynamical interactions, and tidal mass loss from the galactic tidal fields. The Fewbody package is used to integrate directly strong three- and four-body gravitational encounters (Fregeau et al., 2004; Fregeau & Rasio, 2007), which includes post-Newtonian dynamics for BHs (Antognini et al., 2014; Amaro-Seoane & Chen, 2016; Rodriguez et al., 2018a, b). Single and binary star evolution is fully coupled to the stellar dynamics and is computed with COSMIC (Breivik et al., 2020), which is based on SSE (Hurley et al., 2000) and BSE (Hurley et al., 2002).
@@ 2.1 Core-collapse Supernova
BHs and NSs can form in core-collapse supernovae in CMC. The birth mass distribution of compact objects depends on the supernova mechanism and the formation of LMG BHs may be determined by effects including the explosion energy and the development timescale of instabilities (e.g., Fryer etĀ al., 2012; Belczynski etĀ al., 2012; Fryer etĀ al., 2022), or the successfulness of the supernova explosion (e.g., Kochanek, 2014). We follow the prescriptions in Fryer etĀ al. (2012) for the remnant mass and natal kick distributions of core-collapse supernovae and adopt either their ārapid modelā or ādelayed modelā in the simulations. The delayed mechanism launches an explosion in a longer timescale than the rapid mechanism and allows for a broader range of explosion energy and fallback material onto the proto-NS. This mechanism thus produces a continuous range of remnant masses across the LMG (assuming to be between and ), in contrast to the rapid mechanism which leads to a sharp transition between BHs and NSs and produces very few remnants in the mass gap (FigureĀ 4 in Fryer etĀ al. (2012) and also the ārapidā and ādelayedā histograms in FigureĀ 2 below). We adopt a maximum NS mass at following, e.g., Fryer etĀ al. (2012).222The maximum mass of NSs is still highly uncertain and depends on their equation of states (e.g. Ćzel & Freire, 2016), which allows a maximum mass between about and .
We assume that newly born NSs from core-collapse supernovae receive large natal kicks drawn from a Maxwellian distribution with a velocity dispersion (Hobbs etĀ al., 2005). The natal kicks are three-dimensional, with their directions sampled isotropically. On the other hand, the natal kicks of BHs are determined by how much material falls back post-explosion and are calculated by
@@ 2.2 Electron-capture Supernova and White Dwarf Collapse
In addition to core-collapse supernovae, NSs can form from electron-capture supernovae in our simulations (Nomoto, 1984, 1987). Electron-capture can also be triggered in massive white dwarfs (WDs) through accretion onto the WDs or mergers between two WDs (Nomoto & Kondo, 1991; Saio & Nomoto, 1985, 1998, 2004; Shen etĀ al., 2012; Schwab etĀ al., 2015, 2016; Schwab, 2021). We assume that all super-Chandrasekhar mergers where the components are Carbon-Oxygen (CO) or Oxygen-Neon (ONe) WDs collapse to NSs (e.g., Kremer etĀ al., 2023; Ye etĀ al., 2024), thus maximizing the NS formation rate through this channel. For the natal kicks of electron capture supernovae, we adopt the prescription in Kiel etĀ al. (2008) and Kiel & Hurley (2009) and sample the three-dimensional kicks from a Maxwellian distribution with a velocity dispersion . Therefore, many NSs formed in electron capture supernovae are retained in typical GCs with escape velocities of .
@@ 2.3 Collapse of Neutron Stars
To briefly summarize, NS-main sequence TDEs occur when the pericenter distance between the two objects during a close encounter is smaller than the tidal disruption radius of the main-sequence star (e.g., Rees, 1988; Perets etĀ al., 2016):
@@
where , , and are the initial mass and radius of the accretion disk, and the viscous accretion time, respectively (Metzger etĀ al., 2008; Kremer etĀ al., 2022). The exponent is a free parameter determining the amount of mass transported from the outer edge of the accretion disk to the NS (e.g., Blandford & Begelman, 1999), and (e.g., Kumar etĀ al., 2008). We assume that following the results of the hydrodynamic simulations of Kremer etĀ al. (2022). This is a reasonable assumption since the relative velocity of the NSs and main-sequence stars in typical GCs is, in general, much smaller than the stellar escape velocity. We also assume that , day, and equals the NS radius (Kremer etĀ al., 2022). For more details of the TDE treatments in CMC, see Ye etĀ al. (2024).
@@ 3.1 Typical Core-collapse Globular Clusters
We first run 13 core-collapse GC simulations up to a Hubble time (13.8Ā Gyr) with the same initial conditions using CMC to explore the formation of LMG BHs through core-collapse supernovae, merger-induced collapse of WDs, and tidal disruption of stars by NSs. These cluster simulations have initial number of stars , initial virial radius pc, metallicity , and Galactocentric distance kpc. We adopt the Kroupa standard broken power-law (Kroupa, 2001) for the initial mass function between and . The initial binary fraction is assumed to be , representative of the observed binary fraction of old GCs (e.g., Milone etĀ al., 2012).333Studies (e.g., Fregeau etĀ al., 2009) have shown that the overall hard binary fractions of GCs remain roughly constant with time. In addition, because of the large natal kicks for NSs and LMG BHs, varying the initial binary fraction will likely have negligible effects on the formation of binaries with LMG BHs. These simulations will evolve to be ātypicalā core-collapsed clusters similar to NGCĀ 6752 in the Milky Way which shows clear evidence for frequent dynamical interactions of lower-mass compact objects (e.g., single MSPs; for more details, see Ye etĀ al., 2024). Thus clusters similar to NGCĀ 6752 are ideal test grounds for studying the dynamical formation and evolution of LMG BHs. We show the surface brightness profiles and velocity dispersion profiles of all 13 simulations at Gyr in FigureĀ 1 (left panels). We focus on core-collapsed GCs in this study since they produce binaries containing NSs, low-mass BHs, or LMG BHs at higher rates than non-core-collapsed GCs (e.g., Ye etĀ al., 2020).
For these simulations, we vary the supernova prescriptions, the mass scenarios for WDāWD mergers, and the accretion efficiency of NSs tidally disrupting main-sequence stars (SectionĀ 2). These variations will affect the mass distributions of NSs and the number of LMG BHs formed. We also vary the prescriptions for the tidal capture of two WDs, and the stability criterion of binary mass transfer, . This stability criterion will influence the orbital period distribution of the compact object binaries, thereby affecting their cross sections for dynamical encounters (for more details, see Ye etĀ al., 2024). TableĀ 3.2 lists the different prescriptions adopted for each simulation.
To explore the extreme upper limit for the number of LMG BHs and their binaries formed and retained in GCs, we also run a simulation (simulationĀ 14) which truncates the standard Kroupa initial mass function at , resulting in a much smaller number of BHs with masses greater than about . Additionally, we assume that NSs born from core-collapse supernovae receive small natal kicks same as those from electron-capture supernovae. These assumptions allow for a large number of NSs and LMG BHs to be retained in the cluster. Other initial conditions of this simulation are the same as mentioned above.
@@ 3.2 Modeling NGCĀ 1851
In addition to the main grid of models which are intended to be representative of a ātypicalā core-collapsed Milky Way GC, we also run two additional higher mass simulations designed specifically to match the observed properties of NGCĀ 1851. For these models, we assume an initial number of stars and virial radius Ā pc. The metallicity of the simulations is , closely matching the observed metallicity of NGCĀ 1851 from Harris (1996, 2010 edition). We adopt a circular orbit with Galactocentric distance Ā kpc for the cluster following https://people.smp.uq.edu.au/HolgerBaumgardt/globular/, which shows that NGCĀ 1851 is on an eccentric orbit in the Milky Way with pericenter distance Ā kpc and apocenter distance Ā kpc. At distances near the pericenter distance, NGCĀ 1851 may be subject to strong tidal stripping from the Milky Way, and a simple circular orbit at the apocenter distance would not be able to capture the detailed evolution of the cluster in the Milky Wayās potential. However, simulations with Ā kpc match the observed surface brightness profile of NGCĀ 1851 at the outermost part better than those with smaller Galactocentric distances, as we will demonstrate below. This is not surprising since the cluster spends most of its orbital time near the apocenter distance, while only a short time at the pericenter distance. In addition, back scatteringāgravitational interactions with other stars scattering the potential escapers back down to a bound orbitāmay also prevent a significant fraction of objects from escaping from their host clusters even though they have enough energy to do so (e.g., Weatherford etĀ al., 2023). Simulating the detailed evolution of the clusterās tidal boundary is beyond the scope of this paper, and the location of the tidal boundary likely does not have a significant effect on the LMG BHs, which mostly concentrate at the cluster cores due to mass segregation. Other initial cluster properties (e.g., binary fraction, initial mass function, stellar distribution profile) are the same as the simulations in the main grid (simulations 1-13 above). For these two simulations, we again adopt different prescriptions that will affect the mass of the NSs and BHs for comparison (simulationsĀ A and B in TableĀ 3.2). Furthermore, given the observed eccentric orbit of NGCĀ 1851, we include simulationĀ C with Galactocentric distance Ā kpc as a limiting case. All the other initial conditions of this simulation are the same as in simulationĀ B.
@@
3.80 | 24 | 970 | 22 | 1 | 1 | delayed | updated | 0.2 | ā | updated | |
2.23 | 15 | 936 | 20 | 2 | 1 | delayed | updated | 0.2 | ā |
Note. ā From left to right, the columns are the total mass of the cluster at present, the number of BHs retained at present, the number of NSs retained at present, and the number of compact object binaries (NSāWD, NSāNS, and BHāNS) at present. The present-day values for simulations 1-14 are averages between 11 and 13.8Ā Gyr in each cluster simulation, and between 10 and 13Ā Gyr for simulationsĀ A, B, and C.
@@
āSimulationĀ C has the same initial conditions as simulationĀ B except for the Galactocentric distance . SimulationĀ B has Ā kpc while simulationĀ C has Ā kpc.
@@
TableĀ 3.2 summarizes the present-day properties of the 14+2 GC simulations and their prescriptions for forming LMG BHs. The present-day values are averaged between 11 and 13.8Ā Gyr for simulations 1-14, while between 10 and 13Ā Gyr for simulations A-B (the observed age of NGCĀ 1851 is Ā Gyr; VandenBerg etĀ al., 2013). All clusters have reached core-collapsed at late times, which is shown by the small number of BHs retained (e.g., Kremer etĀ al., 2020a, and references therein). In the following sections, we will explore the contributions from various pathways to the formation of binaries and transient sources containing LMG BHs.